3. Suppose a retailer purchased an item with birr 80 and sold for birr 130. Its fixed cost is birr 10,000 and variable cost other than cost of goods sold is 5%.




Compute




A. Develop the cost equation




B. Find Breakeven point in terms of quantity


1
Expert's answer
2022-01-13T09:30:17-0500

A. Let "x=" the number of items sold. Then the total cost is


"C(x)=10000+80(1+0.05)x"

"=10000+84x"

The cost equation is


"C(x)=10000+84x"

B. Total profit at the break-even point is zero.


"10000+84x=130x"

"10000=46x"

"x=218"

"130(218)=28340"

At the break-even point revenue is birr 28340 when 218 items were sold.